# Memory Gym: Partially Observable Challenges to Memory-Based Agents in Endless Episodes

### v1

@@ 3 Memory Baselines
In our experiments, we utilize the popular Deep Reinforcement Learning (DRL) algorithm Proximal Policy Optimization (PPO) (Schulman et al., 2017). By default, PPO lacks a built-in mechanism to enable the ability of memory. To overcome this limitation, either recurrent neural networks (RNN) or self-attention mechanisms (i.e. transformer) are potential solutions. Prior works have extensively demonstrated the effectiveness of leveraging RNNs, such as Long Short-Term Memory (LSTM) (Hochreiter and Schmidhuber, 1997) and Gated Recurrent Unit (GRU) (Cho et al., 2014), within the realm of DRL (Mnih et al., 2016; Espeholt et al., 2018; Huang et al., 2022). This also applies to previous studies employing transformers, like Gated Transformer-XL (GTrXL) (Parisotto et al., 2020) and Hierarchical Chunk Attention Memory (HCAM) (Lampinen et al., 2021). Adopting these memory mechanisms is not trivially plug-and-play. They entail an increase in implementation complexity due to gathering more data, processing sequential data, and extending model interfaces. This may lead to a higher risk of errors and an undesired memory overhead in terms of hardware memory usage. To make these baselines more accessible to the community, we contribute two easy-to-follow baseline implementations of recurrent neural networks and Transformer-XL (TrXL) (Dai et al., 2019) in PPO.It is worth noting that there is a dearth of readily available and accessible transformer-based DRL implementations. Therefore, our realization of TrXL plus PPO is especially valuable in this context. In the following subsections, we provide an overview of the actor-critic model architecture that additionally leverages an auxiliary observation reconstruction loss. Finally, we will delve into the recurrent PPO and transformer-based PPO baselines.
3.1 General Model Architecture
@@
Figure 7: Architecture of the actor-critic model featuring observation encoders and a memory encoder (green) that are shared among the policy, state-value function, and observation reconstruction heads.
@@
Figure 7 presents a broad overview of the actor-critic model architecture. The encoding process begins with visual observations, and optionally, vector observations. To encode visual observations (i.e. pixel observations) of shape $84\times 84\times 3$, we employ the Atari CNN, a convolutional neural network that adheres to the standard topology originally developed for solving Atari games (Mnih et al., 2015). The vector observation is encoded by a multi-layer perceptron (MLP). The outputs of these encoders are concatenated and subsequently embedded into the input space of the memory encoder, which is based on either GRU or Transformer-XL. Once the data is propagated through the memory encoder, it is further processed by each individual head that may contain further hidden layers. The policy head (actor) samples multi-discrete actions as required by Memory Gym, while the value head (critic) approximates the state-value. Additionally, there is a head dedicated to reconstructing the original visual observation, leveraging the principles of self-supervised learning. Technically, the latent output of the memory encoder is fed to a transposed Atari CNN to achieve this reconstruction. The utility of such an auxiliary head is demonstrated by some of our results (Section 4) and prior works by Lampinen et al. (2021) and Hill et al. (2021). During development, we also tried to attach the reconstruction head directly to the encoder, but this led to catastrophic results.
3.2 Loss Function
@@
Our baselines utilize PPO’s clipped surrogate objective (Schulman et al., 2017). Due to the usage of the memory encoder, the to be selected action $a_{t}$ of the policy $\pi_{\theta}$ depends on the current observation $o_{t}$ and the memory encoder’s output that we refer to as hidden state $h_{t}$. $\hat{A}_{t}$ denotes advantage estimates based on generalized advantage estimation (GAE) (Schulman et al., 2016), $\theta$ the parameters of a neural net and $\epsilon$ the clip range.
$L^{C}_{t}(\theta)=-\hat{\mathbb{E}}_{t}[\text{min}(q_{t}(\theta)\hat{A}_{t},\text{clip}(q_{t}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{t})]$ (5)
with $\textnormal{with ratio}~{}q_{t}(\theta)=\frac{\pi_{\theta}(a_{t}|o_{t},h_{t})}{\pi_{\theta\text{old}}(a_{t}|o_{t},h_{t})}$
$t$ depicts the current time step. The value function is optimized using the squared-error loss $L_{t}^{V}(\theta)$. $\mathcal{H}[\pi_{\theta}](o_{t},h_{t})$ denotes an entropy bonus encouraging exploration (Schulman et al., 2017). The auxiliary observation reconstruction loss $L^{R}_{t}$ is based on the binary cross entropy. All three are weighted using the coefficients $c_{1}$, $c_{2}$, and $c_{3}$ and are added to $L_{t}^{C}$ to complete the loss:
$L^{CVHR}_{t}(\theta)=\hat{\mathbb{E}}_{t}[L^{C}_{t}(\theta)+c_{1}L^{V}_{t}(\theta)-c_{2}\mathcal{H}[\pi_{\theta}](o_{t},h_{t})+c_{3}L^{R}_{t}(o_{t},\hat{o}_{t})]$ (6)
$\textnormal{with}~{}L^{V}_{t}(\theta)=(V_{\theta}(o_{t},h_{t})-V_{t}^{target})^{2}$
$\textnormal{and}~{}L^{R}_{t}(o_{t},\hat{o}_{t})=-[\hat{o}_{t}\log(o_{t})+(1-\hat{o}_{t})\log(1-o_{t})]$
3.3 Recurrent Neural Network Baseline
(a) Inference
(b) Optimization
Figure 8: An overview of the distinct forward passes concerning inference and optimization when incorporating an RNN, such as GRU (stacked L times), as a memory encoder in a DRL model. During data sampling (8(a)), the agent interacts with its environment step-by-step, updating the prior hidden state based on encoded features of the current observation. In the optimization phase (8(b)), observations and formerly collected hidden states are divided into sequences. Each sequence is then processed by the RNN with its corresponding initial hidden state, followed by reshaping the resulting data back to its original batch dimensions.
In general, RNNs process sequential data iteratively by propagating information through recurrent connections. When utilizing these in DRL, the forward pass of the model differs between inference (sampling training data) and optimization as shown in Figure 8. During data sampling, the initial hidden state of the recurrent cell is initialized to zeros, and all output hidden states are stored for later optimization. During optimization, the sampled observations, actions, hidden states, and a loss mask are split into sequences based on a configurable sequence length. Sequences shorter than the designated length are zero-padded to optimize computational efficiency, while considering some memory overhead. Each sequence begins with its corresponding initial hidden state, enabling truncated back-propagation through time. Back-propagation through time, introduced by (Werbos, 1990), is an algorithm used to train recurrent neural networks by unfolding the network through time and applying the standard back-propagation algorithm (Rumelhart et al., 1986). Regarding the forward pass during optimization, the model is fed a batch of training data that is reshaped into sequences before being passed to the RNN. Once the data has propagated through the RNN, it is reshaped back to its original batch dimensions. Since only the RNN requires a sequence of data, it is more efficient to feed the non-recurrent components of the neural network the entire batch. Finally, during loss computation, the padded values are excluded using the loss mask, as defined below. $L^{mask}$ is the average over all losses not affected by the padding.
$L^{mask}(\theta)=\frac{\sum_{t}^{T}\left[mask_{t}\times L^{CVHR}_{t}(\theta)\right]}{\sum_{t}^{T}[mask_{t}]}$ (7)
$\textnormal{with}~{}mask_{t}=\begin{cases}0&\text{where padding is used}\\ 1&\text{where no padding is used}\\ \end{cases}$
Our final RNN baseline supports both GRU and LSTM architectures, along with truncated back-propagation through time. For our major experiments we chose GRU over LSTM for two main reasons. Firstly, GRU is computationally less expensive, as demonstrated in the work by Morad et al. (2023). Additionally, it is more GPU memory efficient, as it relies solely on its hidden state without the need for an additional cell state like how it is done in LSTM. Lastly, previous DRL studies showed that GRU slightly outperforms LSTM (Lambrechts et al., 2022; Morad et al., 2023). We also explored LSTM during hyperparameter tuning but observed inferior performance compared to GRU.
3.4 Transformer-XL Baseline
@@
Figure 9: Figure 9(a) illustrates the episodic memory that stores past inputs to all Transformer-XL blocks for every step taken by the agent. Input sequences to the Transformer-XL blocks are retrieved using a sliding memory window. $T$ denotes the episode length. Figure 9(b) depicts the architecture of one Transformer-XL block that is stacked $B$ times. The dashed lines indicate that there is no flow of gradients into the episodic memory, the positional encoding, and the mask.
@@
While RNNs process sequential data by propagating information through recurrent connections, transformers employ self-attention mechanisms to capture global dependencies in parallel across the entire sequence (Vaswani et al., 2017). Next, we detail our transformer baseline that draws conceptual inspiration from TrXL (Dai et al., 2019), GTrXL (Parisotto et al., 2020), and HCAM (Lampinen et al., 2021). The original transformer architecture proposed by Vaswani et al. (2017) is designed as a sequence-to-sequence model with an encoder-decoder structure. However, for the purpose of DRL, we adapt the architecture to a sequence-to-one model, focusing solely on the encoder as done in GTrXL and HCAM. This modification is driven by the underlying Markov Decision Process, which maps the environment’s current state to a single action rather than a sequence of actions.
@@
To begin with, the input data can be conceptualized as a sequence of observations. In the context of visual observations, this is commonly referred to as frame stacking, where consecutive frames are combined to form the input sequence. However, instead of expensively stacking raw frames, we leverage the past outputs of the model’s embedding module (Figure 7), which serves as the input sequence to the first transformer block. The subsequent blocks make use of the past outputs from their preceding blocks. These sequences are stored in the agent’s episodic memory (Figure 9(a)) and are collected during inference based on the agent’s current timestep $t$. To accommodate computational limitations, we employ a sliding memory window approach, where the input sequence is derived from a fixed-length window. By leveraging past activations in the input sequence, the agent’s memory facilitates the ability to attend to events that extend beyond the boundaries of the fixed-length memory window, aligning with the segment-level recurrence mechanism of TrXL (Dai et al., 2019). Note that gradients are not back propagated through the episodic memory.
@@
The architecture of our TrXL encoder block is depicted in Figure 9(b). At each timestep $t$, the block receives an input that is solely based on the current timestep. This input is added to the episodic memory and serves as the query $Q_{t}$ during the execution of multi-head attention (MHA) (Vaswani et al., 2017). The keys $K$ and values $V$ used in MHA are derived from the same data based on self-attention. Specifically, we retrieve $K$ and $V$ by slicing the episodic memory using the bounding indices $i=\max(0,t-\text{{window length}})$ and $t-1$. To ensure the positional information of $K$ and $V$ remains coherent, we add a positional encoding, following the approach from Vaswani et al. (2017). While learning a positional encoding can be effective, we found it to be less sample efficient. To restrict attention to timesteps up to the current timestep $t$, we employ a strictly lower triangular matrix as a mask in MHA. This mask is only necessary when $t$ is smaller than the length of the sliding memory window. The remaining configuration of the block adheres to the findings of Parisotto et al. (2020), which suggest improved performance with pre-layer normalization and the identity map reordering. We encountered vanishing gradient issues when applying layer normalization after MHA and the Position-Wise MLP, which is referred to as post-layer normalization. Although our implementation supports the gating mechanism of GTrXL, it resulted in either lower or equivalent performance while being computationally more expensive (Appendix C).
@@ 4 Experimental Analysis
To benchmark the effectiveness of our just introduced GRU and TrXL baselines, both powered by Proximal Policy Optimization (Schulman et al., 2017), we run empirical experiments on Memory Gym. Contrary to our previous study (Pleines et al., 2023), we exclude HELM (History comprEssion via Language Models) (Paischer et al., 2022a) and its successor, HELMv2 (Paischer et al., 2022b). The decision stems from HELM’s suboptimal wall-time efficiency, which is six times more expensive than our GRU agent, making hyperparameter optimization impractical for our purposes. As we are strongly limited by our compute budget of 25,000 GPU hours, we rather dedicated the resources towards the development of our endless environments and TrXL baseline. Section 5.2 discusses further reasons for not considering more baselines in our experiments. We also conduct preliminary tests using Gated Transformer-XL (GTrXL) (Parisotto et al., 2020) and Long Short-Term Memory (LSTM) (Hochreiter and Schmidhuber, 1997) on the finite environments without any hyperparameter tuning. These models usually underperformed when applied directly, suggesting that out-of-the-box applications might necessitate hyperparameter optimization.
@@ 4.1 Finite Environments: Transformer-XL is often more efficient than GRU
Before assessing GRU and TrXL on the default challenges of the finite environments, the first set of experiments focuses on confirming that the agents trained on Memory Gym indeed require memory. To accomplish this, we train several minor baselines alongside GRU and TrXL on simplified versions of the environments. These baselines include PPO without memory, PPO with frame stacking (4 and 16 grayscale frames), and PPO with relative positional encoding (Vaswani et al., 2017) provided as vector observation.
Results on Mortar Mayhem (Figures 1010(a)-10(c)) show the number of commands that were properly executed during one episode. As mentioned in section 2.1, the agent is not required to memorize the command sequence in Mortar Mayhem Act Grid as it is included in the agent’s vector observation. Under these simplified conditions, PPO without memory is ineffective, while the frame stacking agents achieve nearly either 5 or 8 commands with a slight upward trend. Agents using GRU or TrXL are the most effective, completing all 10 commands, while those with relative positional encoding follow closely, completing 9.5 commands. The positional encoding baseline adds the episode’s current time step to the agent’s observation allowing for temporal awareness within the sequential execution of commands. Once the Clue Task is present (Figure 10(b)), only GRU and TrXL prove to be effective and therefore we conclude that Mortar Mayhem demands memory.
@@
No success is proven by the minor baselines in Searing Spotlights as seen in Figure 10(g). Only in Searing Spotlights, the minor baselines also utilize observation reconstruction. Utilizing the observation reconstruction loss significantly improves the sample efficiency of GRU and TrXL. The reason for this can be found in rare events: the agent, the coin, and the exit are seldomly visible. Hence, the visual encoder takes advantage from the auxiliary learning signal provided by the observation reconstruction. However, both GRU variants, with and without observation reconstruction, exhibit notably higher sample efficiency compared to their respective TrXL counterparts.
When facing the default task of Mortar Mayhem, a different results becomes apparent (Figure 10(c)). TrXL with observation reconstruction solves all 10 commands after training for about 100 million steps, while GRU converges to only about 6 commands. GRU’s performance drops to about 5 commands if the observation reconstruction loss is used. If TrXL trains without this auxiliary loss, its performance drops to about 2 successful commands.
In the case of Mystery Path (Figure 10(f)), we opted not to train agents using observation reconstruction, as the always visible agent is the only entity that requires visual encoding. In this scenario, TrXL achieves the desired success rate in approximately 100 million steps, whereas GRU requires twice as many steps to achieve the same level of performance.

### v3

@@ 4 Transformer-XL Baseline
In our experiments, we employ the widely-recognized Deep Reinforcement Learning (DRL) algorithm Proximal Policy Optimization (PPO) (Schulman et al., 2017). To add memory abilities to PPO, either recurrent neural networks (RNN) or attention mechanisms (i.e. transformer) are potential solutions. Prior works have extensively demonstrated the effectiveness of leveraging RNNs, such as Long Short-Term Memory (LSTM) (Hochreiter and Schmidhuber, 1997) and Gated Recurrent Unit (GRU) (Cho et al., 2014), within the realm of DRL (Mnih et al., 2016; Espeholt et al., 2018; Huang et al., 2022a). This also applies to previous studies employing transformers, as Gated Transformer-XL (GTrXL) (Parisotto et al., 2020) and Hierarchical Chunk Attention Memory (HCAM) (Lampinen et al., 2021). However, integrating these memory-enhancing techniques into DRL algorithms is not straightforward, particularly due to the dearth of readily available transformer-based DRL implementations. To bridge this gap, we contribute an easy-to-follow baseline implementation of Transformer-XL (TrXL) (Dai et al., 2019) in PPO. Before delving into the transformer-based PPO baseline, we describe the actor-critic model architecture and the losses used during training.
4.1 Actor-Critic Model Architecture
@@
Figure 4: Overview of the actor-critic model architecture. The model features observation encoders and a memory encoder (green). The encoders’ parameters are shared among several heads dedicated to the policy, state-value function, observation reconstruction, and ground truth estimation.
@@
Figure 4 presents a broad overview of the actor-critic model architecture. The encoding process begins with visual observations, and optionally, vector observations (i.e. game state information). To encode visual observations (i.e. pixel observations) of shape $84\times 84\times 3$, we employ the Atari CNN (Mnih et al., 2015). The vector observation is encoded by a fully connected layer (FC). The outputs of these encoders are concatenated and subsequently embedded into the input space of the transformer-based memory encoder. Once the data is propagated through the memory encoder, it is further processed by each individual head that may contain further fully connected hidden layers. The policy head (actor) samples multi-discrete actions as required by Memory Gym, while the value head (critic) approximates the state-value.
Optionally, there is a head dedicated to reconstructing the original visual observation to strengthen the learning signal: the latent output of the memory encoder is fed to a transposed Atari CNN. The utility of such an auxiliary head is demonstrated by some of our results (Section 5) and prior works by Lampinen et al. (2021) and Hill et al. (2021). During development, we also tried to attach the reconstruction head directly to the encoder, but this led to catastrophic results. Another optional head establishes a diagnosis tool by estimating ground truth information, which is provided as labels by the environment. Inspired by Baker et al. (2020), the parameters of this head can be optimized post agent training to assess whether distinct information can be queried from the agent’s memory.
4.2 Loss Function Composition
@@
Our baseline utilizes PPO’s clipped surrogate objective (Schulman et al., 2017). Due to leveraging memory, the to be selected action $a_{t}$ of the policy $\pi_{\theta}$ depends on the current observation $o_{t}$ and the memory encoder’s output that we refer to as hidden state $h_{t}$.
$L^{C}_{t}(\theta)=-\mathbb{E}_{t}\left[\text{min}\left(q_{t}(\theta)A_{\pi}^{\text{GAE}}(o_{t},h_{t},a_{t}),\,\text{clip}\left(q_{t}(\theta),1-\epsilon,1+\epsilon\right)A_{\pi}^{\text{GAE}}(o_{t},h_{t},a_{t})\right)\right]$ (1)
with $\textnormal{with ratio}~{}q_{t}(\theta)=\frac{\pi(a_{t}|o_{t},h_{t},\theta)}{\pi(a_{t}|o_{t},h_{t},\theta_{\text{old}})}$
$A_{\pi}^{\text{GAE}}$ denotes advantage estimates based on generalized advantage estimation (GAE) (Schulman et al., 2016), $\theta$ the trainable parameters of a neural net, and $\epsilon$ the clip range. $t$ depicts the current time step. The value function is optimized using the clipped squared-error loss $L_{t}^{V}(\theta)$, which is apparent in the implementation of Schulman et al. (2017):
$L^{VClip}_{t}(\theta)=\left(\text{clip}\left(V(o_{t},h_{t},\theta),V(o_{t},h_{t},\theta_{\text{old}})-\epsilon,V(o_{t},h_{t},\theta_{\text{old}})+\epsilon\right)-\hat{V}\right)^{2}$ (2)
$L_{t}^{V}(\theta)=\max\left[\left(V(o_{t},h_{t},\theta)-\hat{V}\right)^{2},~{}L^{VClip}_{t}(\theta)\right]$ (3)
The auxiliary observation reconstruction is optimized using the binary cross entropy:
$L^{R}_{t}(o_{t},\hat{o}_{t})=-[\hat{o}_{t}\log(o_{t})+(1-\hat{o}_{t})\log(1-o_{t})],$ (4)
and the optional ground truth estimation leverages the squared-error loss, which optimizes the concerned head’s parameters $\phi$:
$L^{Y}_{t}(\hat{y_{t}},y_{t},\phi)=(\hat{y}_{t}-y_{t})^{2}$ (5)
The final combined loss is depicted by $L^{C+V+H+R+Y}_{t}(\theta)$:
$L^{C+V+H+R+Y}_{t}(\theta)=\mathbb{E}_{t}[L^{C}_{t}(\theta)+c_{1}L^{V}_{t}(\theta)-c_{2}\mathcal{H}[\pi_{\theta}](o_{t},h_{t})+c_{3}L^{R}_{t}(o_{t},\hat{o}_{t})+c_{4}L^{Y}_{t}(\theta)]$ (6)
where $\mathcal{H}[\pi_{\theta}](o_{t},h_{t})$ denotes an entropy bonus encouraging exploration (Schulman et al., 2017). $c_{1}$ to $c_{4}$ are coefficients to scale each loss, excluding the clipped surrogate objective.
4.3 Transformer-XL as Memory Mechanism
@@
Figure 5: Figure 5(a) illustrates the episodic memory that stores past inputs to all Transformer-XL blocks for every step taken by the agent. Input sequences to the Transformer-XL blocks are retrieved using a sliding memory window. $T$ denotes the episode length. Figure 5(b) depicts the architecture of one Transformer-XL block that is stacked $L$ times. The dashed lines state that there is no gradient flow into the episodic memory, the absolute positional encoding, and the mask.
@@
When processing sequential data, transformers employ attention mechanisms to capture global dependencies in parallel across the entire sequence (Vaswani et al., 2017), while RNNs iteratively propagate information through recurrent connections. Our transformer baseline draws conceptual inspiration from TrXL (Dai et al., 2019), GTrXL (Parisotto et al., 2020), and HCAM (Lampinen et al., 2021). The original transformer architecture proposed by Vaswani et al. (2017) is designed as a sequence-to-sequence model with an encoder-decoder structure. However, for the purpose of DRL, we adapt the architecture to a sequence-to-one model, focusing solely on the encoder as done in GTrXL and HCAM. This modification is driven by the underlying Markov Decision Process, which maps the environment’s current state to a single action rather than a sequence of actions.
@@
To begin with, it needs to be clarified what kind of sequential data is fed to a transformer, which is typically composed of multiple stacked blocks. Naively, the input data can be conceptualized as a sequence of observations. In the context of visual observations, this is commonly referred to as frame stacking, where consecutive frames are combined. However, instead of expensively stacking raw frames, we leverage the past outputs of the model’s embedding module (Figure 4), which serves as the input sequence to the first transformer block. The subsequent blocks make use of the past outputs from their preceding blocks. These sequences, those elements we refer to as hidden states, are stored in the agent’s episodic memory (Figure 5(a)) and are collected during inference based on the episode’s current timestep $t$. To accommodate computational challenges of potentially endless episodes, we employ a sliding memory window approach, where the input sequence of past hidden states is derived from a fixed-length window. By leveraging formerly computed hidden states in the input sequence, the agent’s memory facilitates the ability to attend to events that extend beyond the boundaries of the fixed-length memory window, aligning with the segment-level recurrence of TrXL (Dai et al., 2019). Note that gradients are not back propagated through the episodic memory. Once propagated through all transformer blocks, a final hidden state is produced to inform the agent’s policy.
@@
The architecture of our TrXL encoder block is depicted in Figure 5(b). At each timestep $t$, the block receives an input that is solely based on the current timestep. This input is added to the episodic memory and serves as the query $Q_{t}$ during the execution of multi-head attention (MHA) (Vaswani et al., 2017). The keys $K$ and values $V$ used in MHA are derived from the same data based on self-attention. Specifically, we retrieve $K$ and $V$ by slicing the episodic memory using the bounding indices $i=\max(0,t-\text{{window length}})$ and $t-1$. To ensure the positional information of $K$ and $V$ remains coherent, we add a positional encoding following the approach from Vaswani et al. (2017). We use the absolute variant as a default as we found a learned one to be effective, but less sample efficient. The underlying positional encoding matrix is scaled to the maximum episode length and not to the length of the window. Therefore, every sequence element of an episode is associate with its distinct position. In the case of endless episodes, the length of the positional encoding is fixed to 2048.
To restrict attention to time steps up to the current timestep $t$, we employ a strictly lower triangular matrix as a mask in MHA. This mask is only necessary when $t$ is smaller than the length of the sliding memory window. The remaining configuration of the block adheres to the findings of Parisotto et al. (2020), which suggest improved performance with pre layer normalization and the identity map reordering. We encountered vanishing gradient issues when applying layer normalization after MHA and the fully connected layer, which is referred to as post layer normalization. Although our implementation supports the gating mechanism of GTrXL, it resulted in either lower or equivalent performance while being computationally more expensive (Appendix C).
@@ 5 Experimental Analysis
We run empirical experiments on Memory Gym using the just introduced TrXL baseline and one based on GRU, both powered by Proximal Policy Optimization (Schulman et al., 2017). Contrary to our previous study (Pleines et al., 2023), we exclude HELM (History comprEssion via Language Models) (Paischer et al., 2022a) and its successor, HELMv2 (Paischer et al., 2022b). The decision stems from HELM’s suboptimal wall-time efficiency, which is six times more expensive than a GRU agent (Pleines et al., 2023), making hyperparameter optimization impractical for our purposes. As we are strongly limited by our compute budget of 25,000 GPU hours, we rather dedicated the resources towards the development of the endless environments and the TrXL baseline. We also conduct preliminary tests using Gated Transformer-XL (GTrXL) (Parisotto et al., 2020) and Long Short-Term Memory (LSTM) (Hochreiter and Schmidhuber, 1997) on the finite environments without any extensive hyperparameter tuning. These models usually underperformed when applied directly, suggesting that out-of-the-box applications might necessitate hyperparameter optimization.
@@ 5.1 Finite Environments: Transformer-XL is often more efficient than GRU
Before assessing GRU and TrXL on the default challenges of the finite environments, the first set of experiments focuses on confirming that the agents trained on Memory Gym indeed require memory. To accomplish this, we train several minor baselines alongside GRU and TrXL on simplified versions of the environments. These baselines include PPO without memory, PPO with frame stacking (4 and 16 grayscale frames), and PPO with absolute positional encoding (Vaswani et al., 2017) provided as vector observation.
Results on Mortar Mayhem (Figures 66(a)-6(c)) show the number of commands that were properly executed during one episode. In Mortar Mayhem Act Grid, where the commands are completely provided as vector observation, PPO without memory is ineffective, while the frame stacking agents achieve nearly either 5 or 8 commands with a slight upward trend. Agents using GRU or TrXL are the most effective, completing all 10 commands, while the one with absolute positional encoding follows closely, completing 9.5 commands. The positional encoding baseline adds the episode’s current time step to the agent’s observation allowing for temporal awareness within the sequential execution of commands. Once the complete task is present (Figure 6(b)), only GRU and TrXL prove to be effective and therefore we conclude that Mortar Mayhem demands memory.
@@
No success is proven by the minor baselines in Searing Spotlights as seen in Figure 6(g). Only in this experiment, the minor baselines also leverage observation reconstruction. Utilizing the observation reconstruction loss significantly improves the sample efficiency of GRU and TrXL. The reason for this can be found in rare events: the agent, the coin, and the exit are seldomly visible. Hence, the visual encoder takes advantage from the auxiliary learning signal provided by the observation reconstruction. However, both GRU variants, with and without observation reconstruction, exhibit notably higher sample efficiency compared to their respective TrXL counterparts.
When facing the default task of Mortar Mayhem, a different result becomes apparent (Figure 6(c)). TrXL with observation reconstruction solves all 10 commands after training for about 100 million steps, while GRU converges to only about 6 commands. GRU’s performance drops to about 5 commands if the observation reconstruction loss is used. If TrXL trains without this auxiliary loss, its performance drops to about only one successful command.
In the case of Mystery Path (Figure 6(f)), we opted not to train agents using observation reconstruction, as the always visible agent is the only entity that requires visual encoding. In this scenario, TrXL achieves the desired success rate in approximately 100 million steps, whereas GRU requires twice as many steps to achieve the same level of performance. If wall-time efficiency is compared, however, TrXL and GRU are roughly on par (Annex D).